OpenStudy (anonymous):

Order the group of quadratic functions from widest to narrowest graph. y=-4x^2, y=-3x^2, y=-5x^2

OpenStudy (neo92):

I think it should be y=-3x^2,y=-4x^2,y=-5x^2

OpenStudy (anonymous):

wider means that if you move some units down in y, which graph will have more distance between it's x's at that y. take at y=-1 for y=-ax^2, calculate x's at y=-1. gives x= +- 1/sqrt(a) or width=2/sqrt(a) So a larger 'a' means narrower

OpenStudy (anonymous):

thanks guys